Suppose you roll a pair of six-sided dice and add their totals. The probability model is given below.
Sum 2 3 4 5 6 7 8 9 10 11 12
Probability 1/36 1/18 1/12 1/9 5/36 1/6 5/36 1/9 1/12 1/18 1/136

(a) What is the probability that the sum of the numbers on your dice is a number other than 12?

EXPLANATION:
The numbers at the top are supposed to be a table. The sum is the top and probability are the bottom. The numbers are aligned (Ex. 2 and 1/36, 3 and 1/18, and so on)

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


The 1/136 should be 1/36

Given table:
Sum23456789101112
Probability1/361/181/121/95/361/65/361/91/121/181/36


This is what it looks like if we transpose the table. This is where any row becomes a column, and vice versa.
SumProbability
21/36
31/18
41/12
51/9
65/36
71/6
85/36
91/9
101/12
111/18
121/36

Side notes:
1/18 = 2/36
1/12 = 3/36
1/9 = 4/36
1/6 = 6/36

According to either table, the probability of rolling a sum of 12 is 1/36.
There's one way to get a sum of 12 (that way being 6+6 = 12) out of 6*6 = 36 outcomes total.

The other 36-1 = 35 outcomes are anything but 12.

Answer: 35/36

Answer by ikleyn(52871)

.

To answer the question, we do not need to know this table.

The only thing that we use in the solution, is the fact that the probability of having the sum 12 is 1/36.

So, the table brings a lot of irrelevant information, which only distracts attention.